OpenStudy (anonymous):

The world speed record on water was set on October 8, 1978 by Ken Warby of Blowering Dam, Australia. If Ken drove his motorboat a distance of 1000 m in 7.045 s, how fast was his boat moving in mi/h.

OpenStudy (shane_b):

Use dimensional analysis...here's a full conversion example: Convert meters to miles:\[\frac{1000\cancel m}{1}*\frac{100\cancel {cm}}{1\cancel m}*\frac{1\cancel {in}}{2.54\cancel {cm}}*\frac{1\cancel {ft}}{12\cancel {in}}*\frac{1mi}{5280\cancel {ft}}=0.6213711922mi\]Convert seconds to hours:\[\frac{7.045\cancel s}{1}*\frac{1\cancel {min}}{60\cancel s}*\frac{1hr}{60\cancel {min}}=0.0019569444hr\]Calculate mph based on the above results:\[\frac{0.6213711922mi}{0.0019569444hr}=317.5211274mi/h\]So that's about 318mi/h.